OpenStudy (anonymous):

find the zeros of the function y=x^2-4x+3 by factoring

OpenStudy (agent0smith):

For x^2+bx+c, find two numbers that multiply to equal c, and add up to b. So find two numbers that multiply to equal +3, and add up to equal -4.

OpenStudy (agent0smith):

Hint: they'll both have to be negative.

OpenStudy (anonymous):

1.5?

OpenStudy (anonymous):

3 and 1

OpenStudy (agent0smith):

They both have to be negative... -3 and -1. (-3)*(-1) = 3 and -3 + (-1) = -4. Now write it as y = (x - 3)(x - 1)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This is confusing

OpenStudy (agent0smith):

From y = (x - 3)(x - 1), to find the zeroes, we let y = 0 0 = (x - 3)(x - 1) Either x-3=0 or x-1=0 x-3 = 0 then x = 3 x-1 = 0 then x = 1